Please explain and show steps. Thank you.

Zambrano Wholesale Corporation maintains its records on a cash basis. At the end of each year the company’s accountant obtains the necessary information to prepare accrual basis financial statements. The following cash flows occurred during the year ended December 31, 2021:

Cash receipts:
From customers $ 705,000
Interest on note 4,480
Loan from a local bank 95,000
Total cash receipts $ 804,480
Cash disbursements:
Purchase of merchandise $ 396,000
Annual insurance payment 6,720
Payment of salaries 216,000
Dividends paid to shareholders 10,000
Annual rent payment 27,000
Total cash disbursements $ 655,720


Selected balance sheet information:

12/31/2020 12/31/2021
Cash $ 28,000 $ 176,760
Accounts receivable 65,000 98,000
Inventory 86,000 65,000
Prepaid insurance 3,100 ?
Prepaid rent 11,600 ?
Interest receivable 3,360 ?
Note receivable 56,000 56,000
Equipment 130,000 130,000
Accumulated depreciation (43,000 ) (56,000 )
Accounts payable (for merchandise) 113,000 128,000
Salaries payable 20,600 25,200
Notes payable 0 95,000
Interest payable 0 ?


Additional information:

  1. On March 31, 2020, Zambrano lent a customer $56,000. Interest at 8% is payable annually on each March 31. Principal is due in 2024.
  2. The annual insurance payment is paid in advance on April 30. The policy period begins on May 1.
  3. On October 31, 2021, Zambrano borrowed $95,000 from a local bank and signed a note promising repayment. Principal and interest at 6% are due on October 31, 2022.
  4. Annual rent on the company’s facilities is paid in advance on June 30. The rental period begins on July 1.


Required:

1. Prepare an accrual basis income statement for 2021 (ignore income taxes).
2. Determine the following balance sheet amounts on December 31, 2021:
Answer #1

solution : 1   Zambrano Wholesale Corporation

INCOME STATEMENT FOR THE YEAR ENDED 31ST DEC.2021

Amount($)
REVENUE: 7,38,000
COST OF SALES (4,97,000)
GROSS PROFIT 2,41,000
LESS: OPERATING EXPENSES
SALARIES 220,600
Rent 25,100
insurance 7580
Deprication 13,000 2,66,280
NON OPERATING INCOME /(LOSS) (25,280)
ADD: INTEREST INCOME 4480
LESS: IN TEREST EXP (950) 3530
NET LOSS (21,750)

WORKING NOTES

REVENUE

CASH SALE =705,000

ADD: CREDIT SALES (INCREASE IN ACCOUNTS RECEIBALE) = 33,000

(98,000 -65000 =33,000)

------------------------------------------------------------------------------------------------------

TOTAL REVENUE = 7,38,000

------------------------------------------------------------------------------------------------------

COST OF SALES

Opening Bal. as on 1Jan21 (inventory)   =86,000

Add: cash purchase =396,000

Add:increase in accounts Payable [128,000 -113000] =15,000

COST OF SALES = 4,97,000

----------------------------------------------------------------------------------------------------------------------------

SALARIES :

SALARIES IN CASH =  216,000

ADD: INCR. SALARIES PAYABLE(25,200 -20,600) = 4600

TOTAL SALARIES = 220,600

----------------------------------------------------------------------------------------------------------------------------Depreciation for the year(56000 -43000) = 13,000

---------------------------------------------------------------------------------------------------------------------------

SOLUTION 2

ACCOUNTS BALANCE

PREPAID INUSRANCE =2240

PREPAID RENT = 13,500

INTEREST RECEIVABLE = 3600

INTEREST PAYABLE = 950

from the information given first i doing journal entries for understanding purpose then we make income statemEnt and calculating remaining accounts balances

s.n Journal Entries DEBIT CREDIT
1 Cash 705,000
Revenue 705,000
2 cash 4480
interest receivable(1 april 20 to 31dec20) 3360
interest income(1jan21 to 31Mar 21) 1120

(Zambrano lent to his customer$ 56,000 , interest receivable@8%annualy each 31 st march,56000 x8% =4480

For the period of 1 april 20 to 31dec20= 4480x9/12 =3360

For the period1jan21 to 31Mar 21= 4480 x 3/12 =1120)

3 interest Receivable(1 april 20 to 31dec20) 3600
interest income(1 april 20 to 31dec20) 3600
4 cash 95,000
notes payable 95,000
5 interest expenses 950
interst payable 950

(loan from bank received on 31st oct 95,000 x6%=5700

interest for 1st nov to 31st dec = 5700x2/12= 950)

6 purchase of merchandise 396,000
cash 396,000
7 insurance 4480
prepaid insurance 2240
cash 6720

(Annual insurance is paid on 30th april &policy beganon 1st may amount paid on 30 april =

insurance for 1st may to 31st dec 21 =6720 x 8/12 =4480

prepaid insurance = 6720 x4/12 = 2240

8 insurance 3100
prepaid insurance(on 31st Dec 20 reversed) 3100
9 salaries 216,000
cash 216,000
10 Dividend 10,000
cash 10,000
11 rent 13,500
prepaid rent 13,500
cash 27,000

(Annual rent is paid in advance on June 30.,rental period is began from 1st july

annual rent paid =27,000

Rent for1 july to 31 dec2021=27000x6/12=13,500

prepaid rent for 1jan22 to 30 june 22=27000 x6/12 =13,500)

12 rent 11600
prepaid rent (as on 31dec 20 reversed) 11600

(please like this ....if you have any query related to this just ask me)